# Break the Checkbox: Challenging Closed-Style Evaluations of Cultural Alignment in LLMs

### v1

@@ 4.2 Hofstede Cultural Dimension
Similar to the WVS results presented in Section 4.1, less constrained settings achieve a higher level of positive correlation with statistical significance in the Hofstede cultural survey. As illustrated in Table 3, $66.67\%$ of the cases of cross-value correlation per country are dominated by the two unconstrained probing methods: forced open-ended and fully unconstrained. Notably, all statistically significant correlations ($p<=0.05$) are exclusively achieved by these two methods. We further evaluate cross-cultural correlations per value across the five models. The findings, as depicted in Figure 4 and Table 11, reveal that while the two closed-style probing methods achieve statistically significant positive correlations in $33.34\%$ of cases across all six dimensions, the two unconstrained settings demonstrate statistically significant results in $58.34\%$ of cases across all dimensions for all models. This further underscores the necessity of employing more flexible, open-ended settings for cultural alignment testing.
One notable observation is that, when prompted in Bengali, the two unconstrained generation settings exhibited suboptimal performance, frequently yielding negative correlations (Table 3) —a trend markedly distinct from the results observed for the other two cultures examined in this study, namely Germany and the USA. Upon closer examination of the responses generated by these unconstrained methods, it becomes evident that the models often fail to provide detailed descriptions or substantive reasoning to support their answers in Bengali. We hypothesize that this underperformance is due to Bengali being a resource-scarce language in the field of natural language processing (NLP), especially when compared to languages like English and German. This disparity arises from limited availability of datasets, underdeveloped NLP tools, and a lack of research focus on Bengali (Joshi et al., 2020). Consequently, the unconstrained generation methods struggle to perform effectively for the Bangladeshi cultural context.
@@ 5 Reconsidering Cultural Alignment
Survey-based closed-style questionnaires are insufficient for accurately evaluating cultural alignment. Anthropological and social science perspectives emphasize the intricate nature of culture, which extends far beyond observable behaviors to encompass the lived experiences of individuals within a society(Geertz, 2017), which cannot be adequately captured through a closed-style multiple-choice question (MCQ) format. Such approach limits language models by preventing them from expressing uncertainty or refusing to answer, both of which can provide deeper insights into societal structures (Urman and Makhortykh, 2023), rather than presenting an artificial sense of alignment through MCQ responses, as discussed in Section 4.1 of our study. A more comprehensive conceptualization of culture views it as an amalgamation of demographic and semantic proxies, each contributing to the complex tapestry of societal norms, values, and behaviors (Adilazuarda et al., 2024), with different evaluation frameworks tailored to assess specific proxies. In this regard, the studies by Naous et al. (2023); Rao et al. (2024) provide valuable examples by developing specialized frameworks to evaluate the adaptability of large language models (LLMs) across various cultural proxies, such as food, regional distinctions, and social values. Based on these insights, we recommend that future research should move beyond the exclusive reliance on closed-style evaluations and adopt more flexible, framework-based methodologies. This approach would allow for a more holistic and accurate assessment of language models’ ability to understand and navigate the complexities of diverse cultural landscapes.
@@
The assertion that language can serve as a reliable proxy for culture is a contentious and potentially flawed approach in academic research. Previous studies have employed language as a representative indicator for cultural alignment and understanding (AlKhamissi et al., 2024; Masoud et al., 2023). However, this methodology presents significant challenges and oversimplifications of the complex relationship between language and culture. The acquisition of a language associated with a particular culture is often a relatively straightforward process that can be accomplished within a relatively short timeframe. In contrast, the internalization and adoption of cultural values is a far more intricate and time-consuming process, often spanning generations or even centuries to become deeply ingrained within a society (Smolicz, 1981). In their study, Havaldar et al. (2023) demonstrate that multilingual models fail to fully capture cultural variations tied to emotions and predominantly reflect the cultural values of Western societies. The ability of a language model to process or generate content in a specific language should not be equated with a comprehensive grasp of the associated cultural values and emotional expressions. This critique is further exemplified by the findings discussed in section 4.2, where the poor performance of unconstrained generation methods in Bengali highlights the practical limitations of NLP models in resource-scarce languages. Together, these observations reinforce the need for both technical advancements in low-resource languages and a conceptual shift beyond relying on language as a sole proxy for culture.
@@
Surveys like WVS and Hofstede can serve as valuable resources of knowledge or fine-tuning data for integrating cultural differences in LLMs , rather than being solely used for evaluating cultural alignment. A notable example of this approach is CultureLLM (Li et al., 2024a), where the authors leverage augmented data derived from WVS survey responses to fine-tune LLMs. Following this fine-tuning process, they evaluate the cultural adaptability of the models using culture-related public datasets in specific languages. This demonstrates the potential of integrating survey data not just for testing but also for improving the cultural sensitivity and adaptability of LLMs.
@@ 6 Conclusion
Closed-style surveys and questionnaires, while effective for collecting human responses, are inadequate for evaluating cultural alignment in large language models (LLMs). Using WVS and Hofstede surveys as a case study, we demonstrate that LLM responses to survey questions exhibit severe instability across varying levels of constraint in the generation settings, as well as with simple changes such as the reordering of choices within the survey questions. We advocate the need for more robust and flexible evaluation frameworks that focus on specific cultural proxies and emphasize localized alignment claims rather than broad global assertions. We believe that this study will open up scopes for developing innovative methodologies for more comprehensive and nuanced evaluations of cultural alignment in LLMs.
@@ Appendix A Background Study
To identify studies that utilize the World Values Survey or Hofstede’s cultural dimensions framework in evaluating large language models (LLMs), we conducted a systematic search across Google Scholar, arXiv, and the ACL Anthology. Our search strategy included the keywords "World Values Survey", "Hofstede", and various synonyms and phrases related to "language models", ensuring comprehensive coverage of relevant literature.
@@ A.1 World Values Survey (WVS)
As the World Values Survey (WVS) is a structured survey in which most questions follow a Likert-scale format, we identified a significant number of studies that have adopted a similar approach when evaluating large language models (LLMs). Most studies adopt a Forced Closed-Style questioning approach, requiring LLMs to generate responses within predefined answer choices to ensure alignment with the original survey design. Zhao et al. (2024) presents WORLDVALUESBENCH, a large-scale benchmark dataset derived from the World Values Survey (WVS) Wave 7, enabling the evaluation of LLMs’ multi-cultural value awareness by predicting human responses to value-based questions based on demographic contexts. Building on value-based assessments, Chiu et al. (2024) examines LLMs’ moral reasoning through DAILYDILEMMAS, a dataset of 1,360 real-life dilemmas, incorporating the World Values Survey (WVS) to analyze cultural value preferences in AI-generated decisions.
Several studies utilize the WVS questionnaire in combination with the Pew Global Attitudes Survey (PEW) to assess cultural values in large language models (LLMs). Durmus et al. (2023) develops a framework for evaluating LLMs’ alignment with global opinions using survey data, incorporating both the World Values Survey (WVS) and the Pew Global Attitudes Survey (PEW) to measure how closely model-generated responses reflect human perspectives across different countries. The study finds that the model’s responses align more closely with opinions from the USA, Canada, Australia, and several European and South American countries compared to other regions. Papadopoulou et al. (2024) replicates and extends prior research on language models’ ability to represent moral norms across cultural contexts, focusing on topics like ’homosexuality’ and ’divorce.’ Using data from the WVS and PEW surveys covering over 40 countries, the analysis reveals that both monolingual and multilingual models exhibit biases and struggle to fully capture the moral complexities of diverse cultures. Ramezani and Xu (2023) investigates whether monolingual English language models can capture moral norms across cultures, focusing on topics like "homosexuality" and "divorce." Using data from the World Values Survey and PEW global surveys, the study finds that pre-trained models struggle to predict cross-cultural moral norms compared to English-specific norms. Fine-tuning on survey data improves predictions for diverse cultures but reduces accuracy for English norms, highlighting challenges in integrating cultural knowledge into moral reasoning.
@@
Lindahl and Saeid (2023) ChatGPT ChatGPT’s values align most closely with developed democracies (e.g., Australia, Great Britain, and Northern Ireland). WVS: MCQ format, sourced from WVS-7.
@@
Cao et al. (2023b) ChatGPT ChatGPT aligns best with American values when prompted with US-based contexts. Hofstede: Multiple-choice (MCQ) format adapted from the Hofstede Culture Survey.
@@ A.2 Hofstede’s Cultural Dimensions
Hofstede’s Cultural Dimensions framework provides a structured approach to analyzing cultural differences across societies. Many studies evaluating large language models (LLMs) have adopted a similar methodology by utilizing multiple-choice and Likert-scale questions derived from Hofstede’s Value Survey Module (VSM). These studies often employ a Forced-Closed-Style questioning method, ensuring that LLM responses align with predefined cultural dimensions, thereby facilitating direct comparison with human survey data. (Kharchenko et al., 2024) investigates the cultural sensitivity of LLMs using Hofstede’s cultural dimensions, analyzing responses to advice-based prompts across 36 countries and multiple languages to assess whether LLMs align with national cultural values. Similarly, Masoud et al. (2023) evaluates LLMs’ alignment with Hofstede’s cultural dimensions using the VSM13 questionnaire, a structured survey with Likert-scale multiple-choice questions. The study finds that GPT-4 demonstrates stronger and more consistent cultural alignment than GPT-3.5 and Llama 2, particularly when adapted to specific personas. Interestingly, despite being primarily trained on English data, GPT-4 aligns more closely with Chinese cultural values while struggling with American and Arab cultural contexts.
@@
Appendix B Evaluation Prompt
@@
Appendix C Anthropological Prompting
Figure 6: Anthropological Prompting.
@@ Appendix D World Values Survey
Table 5: Ten Factors for Two Dimensions of Cross-Cultural Variation for World Values Survey (WVS).
@@ E.5 Unconstrained Questions for Hofstede
Appendix F Additional Findings on Hofstede Cultural Dimension

### v2

@@ 4.2 Hofstede Cultural Dimension
Similar to the WVS results presented in Section 4.1, less constrained settings achieve a higher level of positive correlation with statistical significance in the Hofstede cultural survey. As illustrated in Table 3, $66.67\%$ of the cases of cross-value correlation per country are dominated by the two less constrained probing methods: forced open-ended and fully unconstrained. Notably, all statistically significant correlations ($p<=0.05$) are exclusively achieved by these two methods. We further evaluate cross-cultural correlations per value across the five models. The findings, as depicted in Figure 4 and Table 11, reveal that while the two closed-style probing methods achieve statistically significant positive correlations in $33.34\%$ of cases across all six dimensions, the two unconstrained settings demonstrate statistically significant results in $58.34\%$ of cases across all dimensions for all models.
One notable observation from Table 3 is that, when prompted in Bengali, the two unconstrained generation settings exhibited suboptimal performance, frequently yielding negative correlations-a trend markedly distinct from the results observed for the other two cultures examined in this study, Germany and the USA. Upon closer examination of the responses generated by these unconstrained methods, it becomes evident that the models often fail to provide detailed descriptions or substantive reasoning to support their answers in Bengali. We hypothesize that this underperformance is due to Bengali being a resource-scarce language in the field of natural language processing (NLP), especially when compared to languages like English and German. This disparity arises from limited availability of datasets, underdeveloped NLP tools, and a lack of research focus on Bengali (Joshi et al., 2020). Consequently, the unconstrained generation methods struggle to perform effectively for the Bangladeshi cultural context.
@@ 5 Reconsidering Cultural Alignment
Survey-based closed-style questionnaires are insufficient for accurately evaluating cultural alignment. Anthropological and social science perspectives emphasize the intricate nature of culture, which extends far beyond observable behaviors to encompass the lived experiences of individuals within a society (Geertz, 2017), which cannot be adequately captured through a closed-style multiple-choice question (MCQ) format. Such approach restricts language models by preventing them from expressing uncertainty or refusing to answer—behaviors that can offer valuable insights into societal structures (Urman and Makhortykh, 2023), as discussed in Section 4.1 of our study. Instead, it risks creating an artificial sense of alignment through multiple-choice (MCQ) responses, a phenomenon well-documented in prior literature. A more comprehensive conceptualization of culture views it as an amalgamation of demographic and semantic proxies, each contributing to the complex tapestry of societal norms, values, and behaviors (Adilazuarda et al., 2024), with different evaluation frameworks tailored to assess specific proxies. In this regard, the studies by Naous et al. (2023); Rao et al. (2024) provide valuable examples by developing specialized frameworks to evaluate the adaptability of large language models (LLMs) across various cultural proxies, such as food, regional distinctions, and social values. Based on these insights, we recommend that future research should move beyond the exclusive reliance on closed-style evaluations and adopt more flexible, framework-based methodologies. This approach would allow for a more holistic and accurate assessment of language models’ ability to understand and navigate the complexities of diverse cultural landscapes.
@@
The assertion that language can serve as a reliable proxy for culture is a contentious and potentially flawed approach in academic research. While previous studies have used language as a representative indicator for cultural alignment and understanding (AlKhamissi et al., 2024; Masoud et al., 2023), this methodology oversimplifies the complex relationship between language and culture. The acquisition of a language associated with a particular culture is often a relatively straightforward process that can be accomplished within a relatively short timeframe. In contrast, the internalization and adoption of cultural values is a far more intricate and time-consuming process, often spanning generations or even centuries to become deeply ingrained within a society (Smolicz, 1981). In their study, Havaldar et al. (2023) demonstrate that multilingual models fail to fully capture cultural variations tied to emotions and predominantly reflect the cultural values of Western societies. The ability of a language model to process or generate content in a specific language should not be equated with a comprehensive grasp of the associated cultural values and emotional expressions. This critique is further exemplified by the findings discussed in section 4.2, where the poor performance of unconstrained generation methods in Bengali highlights the practical limitations of NLP models in resource-scarce languages. Together, these observations reinforce the need for both technical advancements in low-resource languages and a conceptual shift beyond relying on language as a sole proxy for culture.
@@
Surveys like WVS and Hofstede can serve as valuable resources of knowledge or fine-tuning data for integrating cultural differences in LLMs , rather than being solely used for evaluating cultural alignment. A notable example of this approach is CultureLLM (Li et al., 2024a), where the authors utilize augmented data derived from WVS survey responses to fine-tune LLMs. After fine-tuning, they evaluate the models’ cultural adaptability by testing them on culture-related public datasets in specific languages, observing measurable improvements. This demonstrates the potential of integrating survey data not only for testing, but also for improving the cultural sensitivity and adaptability of LLMs.
@@ 6 Conclusion
Closed-style surveys and questionnaires, while effective in collecting human responses, are inadequate to evaluate cultural alignment in large language models (LLMs). Using WVS and Hofstede surveys as a case study, we demonstrate that LLM responses to survey questions exhibit severe instability across varying levels of constraint in the generation settings, as well as with simple changes such as the reordering of choices within the survey questions. We emphasize the need for more robust and flexible evaluation frameworks that focus on specific cultural proxies and highlight localized alignment claims rather than broad global assertions. We believe this study will inspire innovative methodologies for more nuanced and comprehensive evaluations of cultural alignment in LLMs.
Limitation
In our study, we specifically focus on the World Values Survey (WVS) and Hofstede’s cultural dimensions framework as case studies to evaluate cultural alignment in large language models (LLMs). While other prominent surveys, such as the Pew Global Attitudes Survey (PEW) have also been utilized in similar contexts, we are unable to conduct experiments across all these datasets due to time and scope constraints. Nevertheless, as discussed in Appendix A, the majority of studies in the field rely predominantly on WVS and Hofstede’s surveys. Many other surveys share a similar multiple-choice format, so the challenges we identify in using WVS and Hofstede’s frameworks are broadly applicable to this general approach to cultural evaluation.
For this work, we focus on three representative cultures from three distinct regions, based on the Inglehart–Welzel Cultural Map. These cultures allow us to derive meaningful and contrasting results regarding the cultural alignment of LLMs. We believe that if we could expand the scope to include a wider variety of cultures from additional regions would yield even deeper and more comprehensive insights. Furthermore, as discussed in Section 4.2, we also hint at the significant issue of low-resource languages in cultural alignment evaluations. The nuances of low-resource languages and their representation in cultural alignment tests have the potential to uncover critical insights and address overlooked gaps in this field. Due to resource constraints, we were unable to explore this dimension further in our current study.
There are virtually limitless variations of probing methods that can be designed to evaluate large language models (LLMs) under varying levels of constraints. In our study, we focus on designing and testing three specific probing methods: forced reverse order, forced open-ended, and fully unconstrained, in addition to the traditional multiple-choice format. These methods are chosen to explore different dimensions of LLM behavior and their ability to align with cultural and contextual nuances. However, we acknowledge that this selection represents only a small subset of the possible probing strategies, leaving significant room for further exploration. Additionally, we evaluate our claims using five state-of-the-art LLMs. While these models were carefully chosen to represent the current advancements in the field, the sheer number of available models presents countless alternative choices. This introduces another limitation to our study, as the results may vary depending on the specific models selected.
@@ Appendix A Background Study
To identify studies that utilize the World Values Survey or Hofstede’s cultural dimensions framework in evaluating large language models (LLMs), we conduct a systematic search across Google Scholar, arXiv, and the ACL Anthology. Our search strategy include the keywords "World Values Survey", "WVS", "Hofstede", and various synonyms and phrases related to "language models", ensuring comprehensive coverage of relevant literature.
@@ A.1 World Values Survey (WVS)
As the World Values Survey (WVS) is a structured survey in which most questions follow either a multiple-choice or Likert-scale format, we identify a significant number of studies that have adopted a similar approach when evaluating large language models (LLMs). Most studies adopt a forced closed-style questioning approach, which requires LLMs to generate responses within predefined answer choices to ensure alignment with the original survey design. Zhao et al. (2024) presents WORLDVALUESBENCH, a large-scale benchmark dataset derived from the World Values Survey (WVS) Wave 7, or WVS 7 in short, enabling the evaluation of LLMs’ multicultural value awareness by predicting human responses to value-based questions based on demographic contexts. Building on value-based assessments, Chiu et al. (2024) examines LLMs’ moral reasoning through DAILYDILEMMAS, a dataset of 1,360 real-life dilemmas, incorporating the World Values Survey (WVS) to analyze cultural value preferences in AI-generated decisions.
Several studies use the WVS questionnaire in combination with the Pew Global Attitudes Survey (PEW) to assess cultural values in large language models (LLMs). Durmus et al. (2023) develops a framework for evaluating LLMs’ alignment with global opinions using survey data, incorporating both the World Values Survey (WVS) and the Pew Global Attitudes Survey (PEW) to measure how closely model-generated responses reflect human perspectives across different countries. The study finds that the model’s responses align more closely with opinions from the USA, Canada, Australia, and several European and South American countries compared to other regions. Papadopoulou et al. (2024) replicates and extends prior research on language models’ ability to represent moral norms across cultural contexts, focusing on topics like ’homosexuality’ and ’divorce.’ Using data from the WVS and PEW surveys covering over 40 countries, the analysis reveals that both monolingual and multilingual models exhibit biases and struggle to fully capture the moral complexities of diverse cultures. In a similar approach, Ramezani and Xu (2023) investigates whether monolingual English language models can capture moral norms across cultures, focusing on topics like "homosexuality" and "divorce." Using data from the World Values Survey and PEW global surveys, the study finds that pre-trained models struggle to predict cross-cultural moral norms compared to English-specific norms. Fine-tuning on survey data improves predictions for diverse cultures but reduces accuracy for English norms, highlighting challenges in integrating cultural knowledge into moral reasoning.
@@
Lindahl and Saeid (2023) ChatGPT WVS: MCQ format, sourced from WVS-7. ChatGPT’s values align most closely with developed democracies (e.g., Australia, Great Britain, and Northern Ireland).
@@
Cao et al. (2023b) ChatGPT Hofstede: Multiple-choice (MCQ) format adapted from the Hofstede Culture Survey. ChatGPT aligns best with American values when prompted with US-based contexts.
@@ A.2 Hofstede’s Cultural Dimensions
Hofstede’s Cultural Dimensions framework provides a structured approach of dividing the factors of a society into six dimensions to analyze cultural differences across societies. Many studies evaluating large language models (LLMs) have adopted a similar methodology by utilizing multiple-choice questions derived from Hofstede’s Value Survey Module (VSM). (Kharchenko et al., 2024) investigates the cultural sensitivity of LLMs using Hofstede’s cultural dimensions, analyzing responses to advice-based prompts across 36 countries and multiple languages to assess whether LLMs align with national cultural values. Similarly, Masoud et al. (2023) evaluates LLMs’ alignment with Hofstede’s cultural dimensions using the VSM13 questionnaire, a structured survey with Likert-scale multiple-choice questions. The study finds that GPT-4 demonstrates stronger and more consistent cultural alignment than GPT-3.5 and Llama 2, particularly when adapted to specific personas. Interestingly, despite being primarily trained on English data, GPT-4 aligns more closely with Chinese cultural values while struggling with American and Arab cultural contexts.
@@
A comprehensive summary of the background study is provided in Table 4 with a concise and structured overview of the key findings and methodologies.
Appendix B Anthropological Prompting
Figure 5: Anthropological Prompting.
Appendix C Evaluation Prompt
@@ Appendix D World Values Survey
Table 5: Ten Factors for Two Dimensions of Cross-Cultural Variation in the World Values Survey (WVS). These questions are used in the Forced Closed-Style probing method for WVS. Each question is preceded by the anthropological prompt (Figure 5) to provide contextual framing at the outset.
@@ D.2 Inglehart–Welzel Cultural Map
D.3 Additional Findings on World Values Survey (WVS)
As outlined in Table 5, the ten indicators are categorized into seven themes: religious values, social values, ethical values, political values, post-materialistic values, happiness and well-being, and social capital. We evaluate our results and compare the four probing methods across these themes using both soft and hard alignment metrics, which are illustrated in Figure groups 7 and 8, respectively.
(a) GPT-4o-Germany-Soft
(b) GPT-4o-Bangladesh-Soft
(c) GPT-4o-USA-Soft
(d) GPT-4-Germany-Soft
(e) GPT-4-Bangladesh-Soft
(f) GPT-4-USA-Soft
(g) Llama 3.3-Germany-Soft
(h) Llama 3.3-Bangladesh-Soft
(i) Llama 3.3-USA-Soft
(j) Mistral Large 2-Germany-Soft
(k) Mistral Large 2-Bangladesh-Soft
(l) Mistral Large 2-USA-Soft
(m) DeepSeek R1-Germany-Soft
(n) DeepSeek R1-Bangladesh-Soft
(o) DeepSeek R1-USA-Soft
Figure 7: Comparison of the four probing methods by themes using the soft alignment metric across all models and countries.
(a) GPT-4o-Germany-Hard
(b) GPT-4o-Bangladesh-Hard
(c) GPT-4o-USA-Hard
(d) GPT-4-Germany-Hard
(e) GPT-4-Bangladesh-Hard
(f) GPT-4-USA-Hard
(g) Llama 3.3-Germany-Hard
(h) Llama 3.3-Bangladesh-Hard
(i) Llama 3.3-USA-Hard
(j) Mistral Large 2-Germany-Hard
(k) Mistral Large 2-Bangladesh-Hard
(l) Mistral Large 2-USA-Hard
(m) DeepSeek R1-Germany-Hard
(n) DeepSeek R1-Bangladesh-Hard
(o) DeepSeek R1-USA-Hard
Figure 8: Comparison of the four probing methods by themes using the hard alignment metric across all models and countries. Some lines representing probing methods are not visible due to overlapping or containing $null$ values.
@@ E.5 Unconstrained Questions for Hofstede
E.6 Additional Findings on Hofstede Cultural Dimension